Rudd Company uses 40,000 micro-chips each year in its production of digital cameras. The of placing an order is $75. The cost of holding one unit of inventory for one year is $8. Currenty Rudd places 20 orders of 2,000 units per order. cost Compute: (a) The annual ordering cost. (2 marks) (2 marks) (2 marks) (2 marks) (b) The annual carrying cost. (c) The total cost of Rudds current inventory policy. (d) The economic order quantity. (e) The order cost and the carrying cost for the EOQ
Answer #1

Ans:

Ans:

a). Annual ordering cost = (40,000 chips / 2000) * 75

= 20 * 75

= $ 1500.

b). Annual carrying cost = (2000 / 2) * 8

= 1000* 8

= $ 8000.

c). Total cost of current inventory policy for the Rudd Company. = 1500 + 8000 = $ 9500.

1). Annual ordering cost

$ 1500

2). Annual carrying cost

$ 8000

3). Total cost of current inventory policy for Rudd Company.

$ 9500.

d) The Economic order quantity

= √(2 * Quantity * Cost per Order / Carrying Cost per Order)

=√(2*40000*75/8)

=√750,000=866.025=866 units

e) Total Costs = Ordering Cost + Carrying Cost

Number of orders=40,000/866=46.1893 orders

Ordering cost = Number of orders*cost of placing an Order

=46.1893*75$ per order=$3464.20

Carrying cost =Quantity Ordered/2*Carrying cost per unit

= 866/2*$8=$3464

Total cost = $6928.20